Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (ending 12 Sep); Correlator hardware (usually 59 BlB pairs) available; some troubling issues with boards; some troubling issues with CC.
  • Key dates:
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
  • Staff Travel:
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
  • Visitors:
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas:EA23 back to the array this week. EA25 out.
    • Receivers: EA24 for first L-band solar capable Rx; several more weeks for this.
      • Perley: RF Switch testing ongoing.
    • Calibration (Perley): Gain compression/expansion ongoing (Hayward).
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Currently 59 BlB pairs ready; some issues from the weekend: board loses communication; specific CCC's begin blocking frames; missing scans (single example to be explored).
  • Computing (Robnett):
    • Notes:

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

  • Mon
    • 1000-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Following showing consistent issues; work this week using phased array.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .Needs close-out testing (Michael-update activation time testing).
  • Correlator: [Clark/Carlson] Delay rocking across bands. (October)
    • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • [Sowinski et al]; Executor update testing
    • Phased array testing (Thursday+)
  • Planetary tracking tests
  • [Owen]: L band deep field study
  • Subarrays: cleanup.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Roy/Dhawan: HI High resolution bandpass testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing.

Commissioning Groups

    • Adv: Low band (Owen):
      • Notes:
    • Adv: Astrometry (Mioduszewski):
      • Notes:Update next week.
  • Calibrator Models
    • Notes: Reduction underway; Ka/S coming very soon; others to follow.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes: ~500 GB data set obtained. Issues noted when filling switched power.
  • Imaging (Ott/Golap):
    • Notes: Takahashi/Ott working on mosaic needs in PST/OPT. Document under review.
  • Algorithm (Bhatnagar, Owen):

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011AConfiguration

Individual reports (updates)

Misc:

Demo Science Status

AOB

-- JosephMcMullin - 2011-09-06
Topic revision: r2 - 2011-09-06, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback